OpenStudy (elite):

What is the y-intercept of the line y = 2x + 6? a0

hartnn (hartnn):

hint : to get y-intercept put x=0 in the equation

hartnn (hartnn):

y = 2(0) + 6 = ... ?

random231 (random231):

^^^^^^

OpenStudy (elite):

uhhm

OpenStudy (elite):

8

OpenStudy (elite):

i guess

hartnn (hartnn):

0+6 is actually = 6 so y intercept is 0,6

OpenStudy (elite):

oh

OpenStudy (elite):

ok thanks

hartnn (hartnn):

welcome ^_^

OpenStudy (elite):

thank you :)